OpenStudy (a1234):

Find all solutions in the interval 0, 2pi sec x + tan x = 1

OpenStudy (freckles):

let u=sec(x) write tan(x) in terms of u to do this you can draw a right triangle using u=sec(x) then write your equation as an algebraic equation and solve for u and then go back and replace u with sec(x) and solve for x

OpenStudy (freckles):



OpenStudy (freckles):

find the opposite side of the angle labeled x using pythagorean theorem

OpenStudy (a1234):

so basically it's 1^2 + b^2 = u^2?

OpenStudy (freckles):

\[b=\pm \sqrt{u^2-1^2} \text{ or } \pm \sqrt{u^2-1}\] since 1^2 is 1...

OpenStudy (freckles):